# Can we soft prompt LLMs for graph learning tasks?

### 15 Feb 2024 v1

@@ Abstract.
Graph plays an important role in representing complex relationships in real-world applications such as social networks, biological data and citation networks. In recent years, Large Language Models (LLMs) have achieved tremendous success in various domains, which makes applying LLMs to graphs particularly appealing. However, directly applying LLMs to graph modalities presents unique challenges due to the discrepancy and mismatch between the graph and text modalities. Hence, to further investigate LLMs’ potential for comprehending graph information, we introduce GraphPrompter, a novel framework designed to align graph information with LLMs via soft prompts. Specifically, GraphPrompter consists of two main components: a graph neural network to encode complex graph information and an LLM that effectively processes textual information. Comprehensive experiments on various benchmark datasets under node classification and link prediction tasks demonstrate the effectiveness of our proposed method. The GraphPrompter framework unveils the substantial capabilities of LLMs as predictors in graph-related tasks, enabling researchers to utilize LLMs across a spectrum of real-world graph scenarios more effectively.
Large Language Models, Graph Neural Networks, Natural Language Processing, Graph Representation Learning
copyright: acmcopyrightjournalyear: 2023copyright: noneconference: The Web Conference 2024; May 13 - 17, 2024; Singapore
@@ 1. Introduction
Large Language Models (LLMs)  (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Radford et al., 2019; Ouyang et al., 2022; Lewkowycz et al., 2022) has demonstrated significant success across various domains, largely attributed to their extensive knowledge memorized during the pretraining phase and their exceptional ability to generalize during the fine-tuning process on diverse textual datasets  (Hoffmann et al., 2022; Webson and Pavlick, 2021; Liang et al., 2022; Carlini et al., 2022). This success has spurred interest in combining graph neural networks (GNNs) with LLMs to enhance their capabilities in understanding and modeling graphs (Yang et al., 2021; Zhang et al., 2022a, b; Ostendorff et al., 2022), including implementing LLMs as encoders to process features within GNNs (Chai et al., 2023; Mavromatis et al., 2023; He et al., 2023; Yu et al., 2023), and employing LLMs as aligners with GNNs to enhance performance (Zhang et al., 2021; Zhao et al., 2022; Wen and Fang, 2023; Brannon et al., 2023).
@@
This hybrid approach is particularly suitable for graphs with textual attributes (i.e., textual graphs), which requires an understanding of both the textual content and the graph structure, such as identifying the subcategory of an academic paper based on its citation network and abstract. We experimentally demonstrate that our framework is capable of prompting LLMs for graph learning tasks on five benchmark datasets under node classification and link prediction tasks. It showcases the potential for significant advancements in the use of LLMs for complex data structures beyond traditional text, opening up new avenues for research and application in the realm of AI assistants capable of intricate graph comprehension. Our main contributions are as follows  [^1] The code is avilable at https://github.com/franciscoliu/graphprompter.:
@@
Figure 1. Illustration of the proposed GraphPrompter for node classification task. The process involves extracting a k-hop subgraph for each node, feeding it into a GNN followed by a projection layer. Simultaneously, the textual attributes associated with each node are processed by the text embedder. The resulting node embedding is then concatenated with text embeddings, serving as a soft promt to guide the LLM for graph learning tasks.
@@ 4. Experiments Result
4.1. Node Classification Task.

### 16 Mar 2024 v2

@@ Abstract.
Graph plays an important role in representing complex relationships in real-world applications such as social networks, biological data and citation networks. In recent years, Large Language Models (LLMs) have achieved tremendous success in various domains, which makes applying LLMs to graphs particularly appealing. However, directly applying LLMs to graph modalities presents unique challenges due to the discrepancy and mismatch between the graph and text modalities. Hence, to further investigate LLMs’ potential for comprehending graph information, we introduce GraphPrompter, a novel framework designed to align graph information with LLMs via soft prompts. Specifically, GraphPrompter consists of two main components: a graph neural network to encode complex graph information and an LLM that effectively processes textual information. Comprehensive experiments on various benchmark datasets under node classification and link prediction tasks demonstrate the effectiveness of our proposed method. The GraphPrompter framework unveils the substantial capabilities of LLMs as predictors in graph-related tasks, enabling researchers to utilize LLMs across a spectrum of real-world graph scenarios more effectively [^1] Code is available at https://github.com/franciscoliu/graphprompter..
Large Language Models; Graph Neural Networks; Natural Language Processing; Graph Representation Learning
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singapore.booktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporeisbn: 979-8-4007-0172-6/24/05doi: 10.1145/3589335.3651476ccs: Computing methodologies Neural networksccs: Information systems Language modelsccs: Computing methodologies Semantic networks
@@ 1. Introduction
Large Language Models (LLMs) have demonstrated significant success across various domains  (Chowdhery et al., 2023; Touvron et al., 2023; Ouyang et al., 2022; Lewkowycz et al., 2022; Wei et al., 2024; Wang et al., 2024), largely attributed to their extensive knowledge memorized during the pretraining phase and their exceptional ability to generalize during the fine-tuning process on diverse textual datasets  (Hoffmann et al., 2022; Liang et al., 2022; Tian et al., 2024a; Carlini et al., 2022; Tan et al., 2024). This success has spurred interest in combining graph neural networks (GNNs) with LLMs to enhance their capabilities in understanding and modeling graphs (Yang et al., 2021; Zhang et al., 2022; Ostendorff et al., 2022; Tian et al., 2024b; He et al., 2024), including implementing LLMs as encoders to process features within GNNs (Chai et al., 2023; Mavromatis et al., 2023; He et al., 2023; Yu et al., 2023), and employing LLMs as aligners with GNNs to enhance performance (Zhao et al., 2022; Wen and Fang, 2023; Brannon et al., 2023).
@@
This hybrid approach is particularly suitable for graphs with textual attributes (i.e., textual graphs), which requires an understanding of both the textual content and the graph structure, such as identifying the subcategory of an academic paper based on its citation network and abstract. We experimentally demonstrate that our framework is capable of prompting LLMs for graph learning tasks on five benchmark datasets under node classification and link prediction tasks. It showcases the potential for significant advancements in the use of LLMs for complex data structures beyond traditional text, opening up new avenues for research and application in the realm of AI assistants capable of intricate graph comprehension. Our main contributions are as follows:
@@
Figure 1. Illustration of the proposed GraphPrompter for node classification task. The process involves extracting a k-hop subgraph for each node, feeding it into a GNN followed by a projection layer. Simultaneously, the textual attributes associated with each node are processed by the text embedder. The resulting node embedding is then concatenated with text embeddings, serving as a soft prompt to guide the LLM for graph learning tasks.
@@ 4. Experiments Result
4.1. Node Classification Task